# Best weighted approximation of some kernels on the real axis

### v1

@@ 3 Main results
$I$ $=$ $\int_{-\infty}^{\infty}|W(t)+Q(t)|^{2}\mathrm{d}t=\int_{-\infty}^{\infty}\Big(W(t)+Q(t)\Big)\Big(\overline{W(t)}+\overline{Q(t)}\Big)\mathrm{d}t.$
@@
$\int_{-\infty}^{\infty}|W(t)|^{2}\mathrm{d}t=\int_{-\infty}^{\infty}W(t)\overline{W(t)}~\mathrm{d}t=\int_{-\infty}^{\infty}Q(t)\overline{Q(t)}~\mathrm{d}t=\int_{-\infty}^{\infty}|Q(t)|^{2}\mathrm{d}t.$
@@
as functions of the complex variable $t$, are regular in the domains $\mathbb{C}_{-}$ and $\mathbb{C}_{+}$, respectively, and for $|t|\to+\infty$ it is ${\mathcal{O}}(|t|^{-l})$. Therefore, based on the Cauchy theorem
$\int_{-\infty}^{\infty}Q(t)\overline{W(t)}~\mathrm{d}t=\int_{-\infty}^{\infty}W(t)\overline{Q(t)}~\mathrm{d}t=0.$
Then
$I=2\int_{-\infty}^{\infty}W(t)\overline{W(t)}\mathrm{d}t=2\int_{-\infty}^{\infty}|W(t)|^{2}\mathrm{d}t.$ (3.19)
@@
$\int_{-\infty}^{\infty}{\textstyle\frac{{\mathrm{d}}t}{({\mathrm{i}}\lambda-t)^{k+1}(-{\mathrm{i}}\lambda-t)^{j+1}}}={\textstyle\frac{\pi}{\lambda}}\cdot{\textstyle\frac{(-1)^{j}{j+k\choose j}}{(2{\mathrm{i}}\lambda)^{j+l}}}.$ (3.20)
@@
$|\tau_{n}({\mathrm{i}\lambda},{\bf\bar{a}})|^{2}=|{\tau}_{n}({-\mathrm{i}\lambda},{\bf a})|^{2}=\prod_{k=1}^{n}[\alpha_{k}^{2}+(\beta_{k}+\lambda)^{2}]=\mu_{n}(\lambda,{\bf a}).$ (3.21)
@@
$I={\textstyle\frac{2\pi}{\lambda\mu_{n}(\lambda,{\bf a})}}\sum_{k=0}^{s}\sum_{l=0}^{s}{\textstyle\frac{(-1)^{l}{l+k\choose l}}{(2{\mathrm{i}}\lambda)^{l+k}}}D_{k}\overline{D}_{l}={\textstyle\frac{4\pi}{(2\lambda)^{2s+3}\mu_{n}(\lambda,{\bf a})}}\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{l+k\choose l}G_{k}\overline{G}_{l},$

### v2

@@ 3 Main results
$I=\int_{-\infty}^{\infty}|W(t)+Q(t)|^{2}\mathrm{d}t=\int_{-\infty}^{\infty}\Big|W(t)\Big|^{2}\mathrm{d}t+\int_{-\infty}^{\infty}\Big|Q(t)\Big|^{2}\mathrm{d}t+2\mathrm{Re}\int_{-\infty}^{\infty}W(t)\overline{Q(t)}\mathrm{d}t.$
@@
$\int_{-\infty}^{\infty}|W(t)|^{2}\mathrm{d}t=\int_{-\infty}^{\infty}|Q(t)|^{2}\mathrm{d}t.$
@@
as functions of the complex variable $t$, are regular in the domains $\overline{\mathbb{C}}_{-}$ and $\overline{\mathbb{C}}_{+}$, respectively, and for $|t|\to+\infty$ it is ${\mathcal{O}}(|t|^{-l})$. Therefore, based on the Cauchy theorem
$\int_{-\infty}^{\infty}Q(t)\overline{W(t)}~\mathrm{d}t=0.$
Therefore
$I=2\int_{-\infty}^{\infty}|W(t)|^{2}\mathrm{d}t.$ (3.19)
@@
$\int_{-\infty}^{\infty}{\textstyle\frac{{\mathrm{d}}t}{({\mathrm{i}}\lambda-t)^{k+1}(-{\mathrm{i}}\lambda-t)^{j+1}}}={\textstyle\frac{\pi}{\lambda}}\cdot{\textstyle\frac{(-1)^{j}{j+k\choose j}}{(2{\mathrm{i}}\lambda)^{j+k}}}.$ (3.20)
@@
$|\tau_{n}({\mathrm{i}\lambda},{\bf\bar{a}})|^{2}=|{\tau}_{n}({-\mathrm{i}\lambda},{\bf a})|^{2}=\frac{|\rho_{n}(-\mathrm{i}\lambda)|}{|\rho_{0}|^{2}}.$ (3.21)
@@
$I={\textstyle\frac{2\pi}{\lambda}}\sum_{k=0}^{s}\sum_{l=0}^{s}{\textstyle\frac{(-1)^{l}{l+k\choose l}}{(2{\mathrm{i}}\lambda)^{l+k}}}D_{k}\overline{D}_{l}={\textstyle\frac{4\pi}{(2\lambda)^{2s+3}\mu_{n}(\lambda,{\bf a})}}\sum_{k=0}^{s}\sum_{l=0}^{s}\textstyle{l+k\choose l}G_{k}\overline{G}_{l},$